OpenStudy (anonymous):

if john made 16 birthday cakes and he sold three of them, and two times of the cakes fell.how many cakes do he have now

OpenStudy (anonymous):

sold cakes 3 cakes fell = 2 times the sold cakes = 2*3 = 6 He had 16 cakes at the beginning sod 3 and 6 fell now how many does he have?

OpenStudy (anonymous):

16- (3 + 6) = 16-9 = 7